OpenStudy (anonymous):

how do you compute this indefinite integral ? ∫(5x^2)/(x^2+1)dx

OpenStudy (shubhamsrg):

1)you can take that 5 out 2)in num,,you have only x^2 ,, write it as (x^2 + 1) -1 3)just a simple simplification and you get 1 - (1/ (x^2 +1)) inside the integral rest should be easier..

OpenStudy (anonymous):

You should break the numerator first I think..

OpenStudy (hba):

do u have options

OpenStudy (anonymous):

\[5\int\limits_{}^{}\frac{x^2 +1 - 1}{x^2 + 1}dx = 5\int\limits_{}^{}dx - 5\int\limits_{}^{}\frac{dx}{x^2 + 1}\]

OpenStudy (anonymous):

For solving later one use the formula: \[\int\limits_{}^{}\frac{dx}{x^2 + a^2} = \frac{1}{a} \tan^{-1} \frac{x}{a} + C\]

OpenStudy (anonymous):

thanks a lot, so when there is x to the power of something in the numerator you always have to simplify the numerator right?

OpenStudy (shubhamsrg):

depends on the denom also..you'll understand better on practicing more..

OpenStudy (anonymous):

how come x+1 didnt get cancel out ?

OpenStudy (anonymous):

No, no it is not necessary..

OpenStudy (anonymous):

x^2 +1 *

OpenStudy (anonymous):

Where what you are talking about??

OpenStudy (anonymous):

See, I show you..